OpenStudy (anonymous):

find the value of t so that the slope of the line through (2,t), and (2,4) is -7

OpenStudy (uri):

Slope formula = (y2-y1) ===== = 7 (x2-x1)

OpenStudy (uri):

-7*

OpenStudy (uri):

4-t == =-7 2-2

OpenStudy (uri):

Solve the equation.

OpenStudy (anonymous):

i got 1.75?